# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Bulk-Surface Interaction

### v2

@@ 1.3  The outline of this paper
The plan of this paper is as follows. In Section 2, we present an equivalent reformulation of system (1.1). Moreover, we compare two simplified variants of our model, one without dynamic boundary conditions and one in a domain with static boundary, with previous models proposed in the literature. We further discuss the benefits of dynamic boundary conditions in our model as well as in Cahn–Hilliard type models in general. Section 3 is then devoted to the derivation of the Navier–Stokes–Cahn–Hilliard model (1.1). We start from scratch by considering local mass balance laws as well as local energy dissipation laws. Eventually, our model derivation is completed via the Langrange multiplier approach.
@@ 2.2  Variants of our model without dynamic boundary conditions
As a further simplification, it is possible to neglect the condition $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$, which also allows us also to omit the surface pressure $q$ as it acts as a Lagrange multiplier for this condition (cf. Subsection 3.3). This simplification was also employed in [50], where a variant of our model with a static boundary was derived. There, neither the condition $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ nor any surface pressure were included in the model. In this way, the system (2.2) further simplifies to
@@
This version can be regarded as a variant of the Abels–Garcke–Grün model (cf. [5]) in an evolving domain with free boundary.
@@ 2.3  The role of dynamic boundary conditions in our model
As already discussed in the literature (see, e.g., [50]), the homogeneous Neumann boundary conditions in (2.3d) are quite restrictive. On the one hand, the condition $\partial_{\mathbf{n}}\phi=0$ on $S_{T}$ means that the contact angle between the diffuse interface separating the fluids and the boundary $\Gamma(\cdot)$ is always exactly ninety degrees. Of course, this is unrealistic in many applications, where the contact angle is expected to change dynamically over the course of time. On the other hand, the condition $\partial_{\mathbf{n}}\mu=0$ on $S_{T}$ implies that no transfer of material between bulk and surface can occur (cf. (1.12)). Therefore, absorption or adsorption processes cannot be described.
These issues are overcome in model (1.1) by means of the bulk-surface convective Cahn–Hilliard subsystem (1.1b)–(1.1f). At least for the non-convective bulk-surface Cahn–Hilliard system (i.e., ${\mathbf{u}}\equiv 0$), it has been observed in several numerical simulations that a transfer of material between bulk and surface occurs if $L\in[0,\infty)$ and that the contact angle between the diffuse interface and the boundary changes dynamically over the course of time (see, e.g., [59, 15, 69, 21, 20]). Unfortunately, the sharp-interface limit of such bulk-surface Cahn–Hilliard models, which would provide an explicit law for the behavior of the contact angle, is still not very well understood.
@@
  • For $K=L=0$, the non-convective bulk-surface Cahn–Hilliard equation was proposed in [39, 52]. Since $L=0$, the chemical potentials are assumed to always remain in a chemical equilibrium, and a rapid transfer of material between bulk and surface is expected to occur.
  • For $K=0$ and $L=\infty$, the non-convective bulk-surface Cahn–Hilliard equation was derived in [63] by an energetic variational approach. In this case, the mass flux between bulk and surface is zero. Consequently, no transfer of material between bulk and surface will occur.
  • The case $K\in(0,\infty)$ and $L=\infty$ was first proposed an analyzed in [58]. It was shown that solutions of this model converge to solutions of the system with $K=0$ and $L=\infty$ as the parameter $K$ is sent to zero.
  • For $K=0$ and $L\in(0,\infty)$, the non-convective bulk-surface Cahn–Hilliard equation was proposed in [59] to interpolate between the models associated with $K=0$, $L=0$ and K=0{K=0}, $L=\infty$}. In this case, a transfer of material between bulk and surface will occur, and the number $L^{-1}$ is related to the rate of absorption (or adsorption) of bulk material by the boundary (cf. [59]).
  • In [61, 62, 51], the convective bulk-surface Cahn–Hilliard equation with a given velocity field was studied for all cases $K,L\in[0,\infty]$ in a unified framework. Of course, these results also include the non-convective case by simply choosing ${\mathbf{u}}\equiv 0$ for the prescribed velocity field.
@@ 2.4  Variants of our model with a static boundary
where $Q_{T}$ and $S_{T}$ are to be interpreted as $Q_{T}=\Omega\times(0,T)$ and $S_{T}=\Gamma\times(0,T)$, respectively. This model can be understood as a variant of the Navier–Stokes–Cahn–Hilliard model derived in [50], which further includes the condition $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$ along with the surface pressure $q$ acting as a Lagrange multiplier for this condition. We point out that it becomes clear by the model derivation (see Section 3) that the boundary condition (1.1g) drops out when the condition ${\mathbf{u}}\cdot{\mathbf{n}}=0$ on $S_{T}$ is demanded.
If we further neglect the relation $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$ and set $q\equiv 0$, we arrive exactly at the model that was derived in [50]. A Cahn–Hilliard–Brinkman variant of this model, in which the Navier–Stokes equation is replaced by a Brinkman equation to describe flows through porous media, was investigated in [30].
@@ 3  Model derivation
In this section we will derive the Navier–Stokes–Cahn–Hilliard system with dynamic boundary conditions (1.1).
We consider an evolving domain $\{\Omega(t)\}_{t\in[0,T]}$ in $\mathbb{R}^{d}$ of class $C^{2}$. For every $t\in[0,T]$, let $\Gamma(t)\coloneqq\partial\Omega(t)$. This means that $\{\Gamma(t)\}_{t\in[0,T]}$ is a closed, oriented evolving hypersurface in $\mathbb{R}^{d}$ of class $C^{2}$. For brevity, we use the notation
$Q_{T}\coloneqq\bigcup_{0<t<T}\{t\}\times\Omega(t)\quad\text{and}\quad S_{T}\coloneqq\bigcup_{0<t<T}\{t\}\times\Gamma(t).$
We further assume that the evolving domain $Q_{T}$ is transported by a vector field ${\mathbf{u}}\in C^{2}(\overline{Q_{T}})$, which is supposed to be the volume-averaged velocity field of the two fluids that are considered. Therefore, $S_{T}$ is transported by the vector field ${\mathbf{u}}|_{S_{T}}\in C^{2}(S_{T})$. This means that
$\Omega(s)=\eta_{s}\big(t,\Omega(t)\big)\quad\text{and}\quad\Gamma(s)=\eta_{s}\big(t,\Gamma(t)\big)\quad\text{for all $s,t\in[0,T]$,}$
where for any $(t,{\mathbf{x}})\in\overline{Q_{T}}$, the flow map $0,T]\ni s\mapsto\eta_{s}(t,{\mathbf{x}})\in\overline{\Omega(s)}$ is the unique solution of the ordinary differential equation
$\frac{\mathrm{d}}{\mathrm{d}s}\eta_{s}(t,{\mathbf{x}})={\mathbf{u}}\big(s,\eta_{s}(t,{\mathbf{x}})\big),\quad s\in[0,T],$
which satisfies the initial condition $\eta_{s}(t,{\mathbf{x}})|_{s=t}={\mathbf{x}}$. Since $S_{T}$ is orientable, there exists an evolving unit normal vector field ${\mathbf{n}}\in C(S_{T})$ and the normal velocity $\mathcal{V}$ of $S_{T}$ is given by
@@
For more details about the theory of evolving domains and hypersurfaces as well as corresponding calculus and transport theorems, we refer to [13, 17].
@@ 3.1  Considerations based on local mass balance laws
Here, $\widehat{\mathbf{J}}_{i}$ are the mass fluxes corresponding to the motion of the two materials in the bulk. In our model, we also want to account for a transfer of material between bulk and surface. This is, for instance, needed in order to describe absorption processes. Therefore, we also consider functions $\sigma_{i}:S_{T}\to\mathbb{R}$, $i=1,2$, which represent the mass densities of the substrates on the free boundary. In general, the materials on the surface might differ from those in the bulk. This is the case, for example, if the materials are transformed by chemical reactions occurring at the boundary. Therefore, we interpret the densities $\sigma_{i}$ as independent functions that are not necessarily related to the traces of the densities $\rho_{i}$ (up to a constant carrying physical units), respectively. We assume that the the following mass balance law holds on the free boundary:
@@
We now introduce $\phi_{i}:Q_{T}\to\mathbb{R}$ and $\psi_{i}:S_{T}\to\mathbb{R}$, $i=1,2$, to denote the volume fractions of the two fluids in the bulk and the two substrates on the boundary, respectively. Recalling that each component has a constant density $\tilde{\rho}_{i}$, $\tilde{\sigma}$, $i=1,2$, the volume fractions can be identified as
@@
which is frame indifferent. Hence, (3.23) can be rewritten as
@@
Multiplying the equations (3.15) each by ${1}/{\tilde{\rho}_{i}}$, $i=1,2$, and subtracting the resulting equations, we deduce that
$\partial_{t}\phi+\textnormal{div}\,(\phi{\mathbf{u}})+\textnormal{div}\,\,\mathbf{J}_{\phi}=0\quad\text{in $Q_{T}$.}$ (3.31)
@@
$\textnormal{div}_{\Gamma}(\tilde{\mathbf{K}}_{1}+\tilde{\mathbf{K}}_{2})$ $=-\partial_{t}^{\circ}(\psi_{1}+\psi_{2})-\textnormal{div}_{\Gamma}\big((\psi_{1}+\psi_{2}){\mathbf{u}}_{\tau}\big)+\big(\tilde{\mathbf{J}}_{\Gamma,1}+\tilde{\mathbf{J}}_{\Gamma,2}\big)\cdot{\mathbf{n}}=0$ (3.34)
@@
This means that the fluxes $\mathbf{J}$ and $\frac{1}{2}(\tilde{\rho}_{2}-\tilde{\rho}_{1})\mathbf{J}_{\phi}$ can merely differ by an additive divergence-free function. As in [5, 50], we thus assume that
@@
In the same way, we derive the identity
@@
3.2  Local energy dissipation laws
As in [5] and [50], we consider the following energy density in the bulk:
@@ 3.2  Local energy dissipation laws
As in [50], we additionally introduce the surface energy density
@@
$g(\phi,\psi,\nabla_{\Gamma}\psi)=\frac{\delta}{2}\left|\nabla_{\Gamma}\psi\right|^{2}+\frac{1}{\delta}G(\psi)+\mathds{h}(K)(\alpha\psi-\phi)^{2},\quad\text{$K\in[0,\infty]$}$ (3.42)
@@
In this inequality, the vector fields $\mathbf{J}_{e}$ and $\mathbf{K}_{e}$ are energy fluxes that will be specified later, ${\mathbf{n}}_{\partial V(t)}$ is the outer unit normal vector field on $\partial V(t)$, and ${\mathbf{n}}_{\partial(\partial V(t)\cap\Gamma(t))}$ is the unit conormal vector field on the relative boundary $\partial\big(\partial V(t)\cap\Gamma(t)\big)$ within the submanifold $\Gamma(t)$. As we consider a thermodynamically closed system, there is no exchange of energy over the boundary $\Gamma(t)$ and thus, the domain of the first integral in the second line is just $\partial V(t)\cap\Omega(t)$ instead of $\partial V(t)$. Applying Gauß’s divergence theorem for not necessarily tangential vector fields (see, e.g., [13, Proposition 2.5.1]) on both integrals in the second line, we reformulate (3.43) as
@@
Here, $H=-\frac{1}{d}\textnormal{div}_{\Gamma}{\mathbf{n}}$ is the mean curvature of $\Gamma(t)$, and we point out that ${\mathbf{n}}_{\partial V(t)}={\mathbf{n}}(t)$ on $\Gamma(t)$. Applying the Reynolds transport theorem for evolving domains (see, e.g., [13, Theorem 2.11.1]), we deduce the identity
@@
Similarly, recalling that $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ (see (3.27)), we use the transport theorem for orientable evolving hypersurfaces (see, e.g., [13, Theorem 2.10.1]) to obtain
@@
3.3  Completion of the model derivation by the Lagrange multiplier approach
To complete the model derivation, we employ the Lagrange multiplier approach (see [64]). Therefore, we introduce the Lagrange multipliers $\mu$, $\theta$ and $q$ that need to be appropriately adjusted in the following approach. In the final model, $\mu$ will represent the bulk chemical potential, $\theta$ will represent the surface chemical potential and $q$ will represent the surface pressure. Invoking the identities (3.31) and (3.33), the local energy dissipation laws (3.48) and (3.49) can be expressed as

### v3

@@ 1.3  The outline of this paper
The plan of this paper is as follows. In Section 2, we present an equivalent reformulation of system (1.1). Moreover, we compare two simplified variants of our model, one without dynamic boundary conditions and one in a domain with static boundary, with previous models proposed in the literature. We further discuss the benefits of dynamic boundary conditions in our model as well as in Cahn–Hilliard type models in general. Section 3 is then devoted to the derivation of the Navier–Stokes–Cahn–Hilliard model (1.1). Starting from the local mass balance laws, we derive the model by two different approaches: the Lagrange Multiplier Approach and, in the case of matched densities and no mass flux between bulk and surface, the Energetic Variational Approach.
@@ 2.2  Variants of our model without dynamic boundary conditions
As a further simplification, it is possible to neglect the condition $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$, which also allows us also to omit the surface pressure $q$ as it acts as a Lagrange multiplier for this condition (cf. Subsection 3.2.2). This simplification was also employed in [Giorgini2023], where a variant of our model with a static boundary was derived. There, neither the condition $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ nor any surface pressure were included in the model. In this way, the system (2.2) further simplifies to
@@
This version can be regarded as a variant of the Abels–Garcke–Grün model (cf. [Abels2012]) in an evolving domain with free boundary.
@@ 2.3  The role of dynamic boundary conditions in our model
As already discussed in the literature (see, e.g., [Giorgini2023]), the homogeneous Neumann boundary conditions in (2.3d) are quite restrictive. On the one hand, the condition $\partial_{\mathbf{n}}\phi=0$ on $S_{T}$ means that the contact angle between the diffuse interface separating fluids and the boundary $\Gamma(\cdot)$ is always exactly ninety degrees. Of course, this is unrealistic in many applications, where the contact angle is expected to change dynamically over the course of time. On the other hand, the condition $\partial_{\mathbf{n}}\mu=0$ on $S_{T}$ implies that no transfer of material between bulk and surface can occur (cf. (1.12)). Therefore, absorption or adsorption processes cannot be described.
These issues are overcome in model (1.1) by means of the bulk-surface convective Cahn–Hilliard subsystem (1.1b)–(1.1f). At least for the non-convective bulk-surface Cahn–Hilliard system (i.e., ${\mathbf{u}}\equiv 0$), it has been observed in several numerical simulations that a transfer of material between bulk and surface occurs if $L\in[0,\infty)$ and that the contact angle between the diffuse interface and the boundary changes dynamically over the course of time (see, e.g., [Knopf2021a, Bao2021a, Metzger2023, Bullerjahn2024, Bullerjahn2025]). Unfortunately, the sharp-interface limit of such bulk-surface Cahn–Hilliard models, which would provide an explicit law for the behavior of the contact angle, is still not very well understood. In [Metzger2025, Zhang2023], the relation between the Cahn–Hilliard equation with dynamic boundary conditions and a generalized Young’s formula for the contact angle was discussed at least on a heuristic level.
@@
  • For $K=L=0$, the non-convective bulk-surface Cahn–Hilliard equation was proposed in [Gal2006, Goldstein2011]. Since $L=0$, the chemical potentials are assumed to always remain in a chemical equilibrium, and a rapid transfer of material between bulk and surface is expected to occur.
  • For $K=0$ and $L=\infty$, the non-convective bulk-surface Cahn–Hilliard equation was derived in [Liu2019] by the Energetic Variational Approach. In this case, the mass flux between bulk and surface is zero. Consequently, no transfer of material between bulk and surface will occur.
  • The case $K\in(0,\infty)$ and $L=\infty$ was first proposed an analyzed in [Knopf2020]. It was shown that solutions of this model converge to solutions of the system with $K=0$ and $L=\infty$ as the parameter $K$ is sent to zero.
  • For $K=0$ and $L\in(0,\infty)$, the non-convective bulk-surface Cahn–Hilliard equation was proposed in [Knopf2021a] to interpolate between the models associated with $K=0$, $L=0$ and K=0{K=0}, $L=\infty$}. In this case, a transfer of material between bulk and surface will occur, and the number $L^{-1}$ is related to the rate of absorption (or adsorption) of bulk material by the boundary (cf. [Knopf2021a]).
  • In [Knopf2024, Knopf2024a, Giorgini2025], the convective bulk-surface Cahn–Hilliard equation with a given velocity field was studied for all cases $K,L\in[0,\infty]$ in a unified framework. Of course, these results also include the non-convective case by simply choosing ${\mathbf{u}}\equiv 0$ for the prescribed velocity field.
@@ 2.4  Variants of our model with a static boundary
where $Q_{T}$ and $S_{T}$ are to be interpreted as $Q_{T}=\Omega\times(0,T)$ and $S_{T}=\Gamma\times(0,T)$, respectively. This model can be understood as a variant of the Navier–Stokes–Cahn–Hilliard model derived in [Giorgini2023], which further includes the condition $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$ along with the surface pressure $q$ acting as a Lagrange multiplier for this condition. We point out that it becomes clear by the model derivation (see Section 3) that the boundary condition (1.1h) drops out when the condition ${\mathbf{u}}\cdot{\mathbf{n}}=0$ on $S_{T}$ is demanded.
If we further neglect the relation $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$ and set $q\equiv 0$, we arrive exactly at the model that was derived in [Giorgini2023]. A Cahn–Hilliard–Brinkman variant of this model, in which the Navier–Stokes equation is replaced by a Brinkman equation to describe flows through porous media, was investigated in [Colli2024].
@@ 3  Model derivation
In this section we will derive the Navier–Stokes–Cahn–Hilliard system with dynamic boundary conditions (1.1). Let $\{\Omega(t)\}_{t\in[0,T]}$, $\{\Gamma(t)\}_{t\in[0,T]}$, $Q_{T}$, and $S_{T}$ be as introduced in Subsection 1.1. Recall that the evolving domain $Q_{T}$ is transported by a vector field $\mathbf{u}$}, which is supposed to be the volume-averaged velocity field of the two fluids that are considered. Therefore, $S_{T}$ is transported by the vector field ${\mathbf{u}}|_{S_{T}}$. Moreover, since $S_{T}$ is orientable, there exists an evolving unit normal vector field ${\mathbf{n}}\in C(S_{T})$ and the normal velocity $\mathcal{V}$ of $S_{T}$ is given by
@@ 3.1  Considerations based on local mass balance laws
Here, $\widehat{\mathbf{J}}_{i}$ are the mass fluxes corresponding to the motion of the two materials in the bulk. In our model, we also want to account for a transfer of material between bulk and surface. This is, for instance, needed in order to describe absorption processes. Therefore, we also consider functions $\sigma_{i}:S_{T}\to\mathbb{R}$, $i=1,2$, which represent the mass densities of the substrates on the free boundary. In general, the materials on the surface might differ from those in the bulk. This is the case, for example, if the materials are transformed by chemical reactions occurring at the boundary. Therefore, we interpret the densities $\sigma_{i}$ as independent functions that are not necessarily related to the traces of the densities $\rho_{i}$ (up to a constant carrying physical units), respectively. We assume that the following mass balance law holds on the free boundary:
@@
We now introduce $\phi_{i}:Q_{T}\to\mathbb{R}$ and $\psi_{i}:S_{T}\to\mathbb{R}$, $i=1,2$, to denote the volume fractions of the two fluids in the bulk and the two substrates on the boundary, respectively. Assuming that the bulk and surface materials each have a constant density $\tilde{\rho}_{i}$, $\tilde{\sigma}_{i}$, $i=1,2$, the volume fractions can be identified as
@@
Hence, (3.23) can be rewritten as
@@
Multiplying the equations (3.15) each by ${1}/{\tilde{\rho}_{i}}$, $i=1,2$, and subtracting the resulting equations, and recalling (3.26), we deduce that
$\partial_{t}^{\bullet}\phi+\textnormal{div}\,\,\mathbf{J}_{\phi}=\partial_{t}\phi+\textnormal{div}\,(\phi{\mathbf{u}})+\textnormal{div}\,\,\mathbf{J}_{\phi}=0\quad\text{in $Q_{T}$.}$ (3.31)
@@
[…]
@@
This means that there exists a divergence-free function $\mathbf{J}_{0}$ such that
$\mathbf{J}=\frac{\tilde{\rho}_{2}-\tilde{\rho}_{1}}{2}\mathbf{J}_{\phi}+\mathbf{J}_{0}\quad\text{in $Q_{T}$.}$ (3.36)
It is a common assumption in mixture theory that the relation between $\mathbf{J}$ and $\mathbf{J}_{\phi}$ is linear, that is, $\mathbf{J}_{0}=\mathbf{0}$ in $Q_{T}$. This means that
@@
Alternatively, this linear relation can also be shown by assuming the final model to satisfy the global energy dissipation law (1.10). Using the relation (3.36) would produce an additional summand $\textnormal{div}\,({\mathbf{u}}\otimes\mathbf{J}_{0})$ on the left-hand side of the Navier–Stokes equation (1.1a). However, the energy dissipation law (1.10) can only be satisfied if $\mathbf{J}_{0}$ is zero. Hence, in the following, we readily assume that $\mathbf{J}_{0}=\mathbf{0}$ in $Q_{T}$.
Proceeding similarly to (3.35), we further derive the identity
@@
3.2  Option A: Model derivation via the Lagrange Multiplier Approach
One possibility to complete the model derivation is the Lagrange Multiplier Approach (see [Liu2002]). It relies on local energy dissipation laws, the introduction of Lagrange multipliers, and constitutive assumptions to identify the unkown flux terms.
3.2.1 Local energy dissipation laws
As in [Abels2012] and [Giorgini2023], we consider the following energy density in the bulk:
@@ 3.2.1 Local energy dissipation laws
As in [Giorgini2023], we additionally introduce the surface energy density
@@
$g(\phi,\psi,\nabla_{\Gamma}\psi)=\frac{\delta}{2}\left|\nabla_{\Gamma}\psi\right|^{2}+\frac{1}{\delta}G(\psi)+\frac{1}{2}\mathds{h}(K)(\alpha\psi-\phi)^{2},\quad\text{$K\in[0,\infty]$}$ (3.43)
@@
In this inequality, the vector fields $\mathbf{J}_{e}$ and $\mathbf{K}_{e}$ are energy fluxes that will be specified later, ${\mathbf{n}}_{\partial V(t)}$ is the outer unit normal vector field on $\partial V(t)$, and ${\mathbf{n}}_{\partial(\partial V(t)\cap\Gamma(t))}$ is the unit conormal vector field on the relative boundary $\partial\big(\partial V(t)\cap\Gamma(t)\big)$ within the submanifold $\Gamma(t)$. As we consider a thermodynamically closed system, there is no exchange of energy over the boundary $\Gamma(t)$ and thus, the domain of the first integral in the second line is just $\partial V(t)\cap\Omega(t)$ instead of $\partial V(t)$. Applying Gauß’s divergence theorem for not necessarily tangential vector fields (see, e.g., [BDGP-book, Proposition 2.5.1]) on both integrals in the second line, we reformulate (3.44) as
@@
Here, $H=-\frac{1}{d}\textnormal{div}_{\Gamma}{\mathbf{n}}$ is the mean curvature of $\Gamma(t)$, and we point out that ${\mathbf{n}}_{\partial V(t)}={\mathbf{n}}(t)$ on $\Gamma(t)$. Applying the Reynolds transport theorem for evolving domains (see, e.g., [BDGP-book, Theorem 2.11.1]), we deduce the identity
@@
Similarly, recalling that $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ (see (3.27)), we use the transport theorem for orientable evolving hypersurfaces (see, e.g., [BDGP-book, Theorem 2.10.1]) to obtain
@@
3.2.2 The Lagrange Multiplier Approach
Now, we introduce the Lagrange multipliers $\mu$, $\theta$ and $q$ that need to be appropriately adjusted in the following approach. In the final model, $\mu$ will represent the bulk chemical potential, $\theta$ will represent the surface chemical potential and $q$ will represent the surface pressure. Invoking (3.26), (3.27), (3.31), and (3.33), the local energy dissipation laws (3.49) and (3.50) can be expressed as